Can someone help e solve this problem?
I need to solve: 10x - 1/4 = 9x + 3/8
Thank you,
sher

Answer by checkley71(8403)

10X-1/4=9X+3/8
10X-9X=3/8+1/4
X=3/8+2/8
X=5/8
PROOF
10*5/8-1/4=9*5/8+3/8
50/8-1/4=45/8+3/8
50/8-2/8=(45+3)/8
(50-2)/8=48/8
48/8=48/8
6=6